Weekly Inspection – April Rain

Observations:

Well the bad weather has continued this week, lots of rain and temperatures only reaching 12°C. The rain meant that I put off doing the last stage of the queen rearing but only by a day. Noticeable that the bees haven’t been able to forage much and have been eating through their stores this week. Transferred sealed queen cells into the two apideas. Replaced one of the sealed queen cells with a spare laying queen I was given from two colonies being united (not my bees). I set up a nuc with brood and bees from Aqua and Cyan – dusted with icing sugar to prevent fighting and put the queen into a Butler cage between two frames. So I’ll see how she gets on over the next few weeks.

On Thursday, 19th whilst doing the transfer of queen cells in the Plum colony I examined the Maroon colony which had shown signs of swarming last week. There were more new charged queen cells so I took out the queen and four frames of brood and stores and put into a nuc and brought it back to my home apiary. I will let the parent colony raise a new queen.

So I currently have 2 new nucleus colonies(Green and Lime), each with laying queens.

Took out the varroa slides at the Mole apiary and only found a total of 3 mites, so nothing to worry about on this front.

Flowering: Pear and horse chesnut both coming into bloom.

Hive Record
Colony
Q
BIAS
Brood
Stores
CommentNext Inspection
AquaYY9 frames in bottom box and two in top box. Took out two for the new nuc.About 6Kgs in the top box.Looks to be in good shape.Check for swarming signs.
BlueYY3 frames of broodCouple of framesDoesn't appear to be expanding as it's not replaced the extra frame of donated brood.Consider uniting with the nuc (Green) brought from Mole apiary
CyanYY9 frames of brood. 1 removed with additional frames of stores for new nuc.Still not making in roads on the superLooks to be doing OK.Check if frames need re-arranging, after withdrawal of 2 today
MaroonNY5 frames of brood (minus 3 for the nuc)About 4Kgs - down from a last weekCurrently 7 charged queen cellsCheck mid-week for sealed queen cells
OrangeYY9 frames of broodAbout 1 Kg in superLooks to be in good orderCheck for space in brood nest
PlumYY5 frames of broodAbout 3 framesLooks to be plenty of bees, brood nest has shrunk. Certainly seems to have stopped swarming beahaviour. Depends on queen raiser above
LimeYY3 frames of brood1/2 frame + syrup feederNew nuc set-up from home apiary brood and beesFeed during mid-week. Check for eggs
GreenYY3 frames of broodContact feeder - taken about 1 litre of syrup so farQueen looks to be layingCheck for stores. Possible unite with Blue nuc